# Reinforced Latent Reasoning for LLM-based Recommendation

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive reasoning capabilities in complex problem-solving tasks, sparking growing interest in their application to preference reasoning in recommendation systems. Existing methods typically rely on fine-tuning with explicit chain-of-thought (CoT) data. However, these methods face significant practical limitations due to (1) the difficulty of obtaining high-quality CoT data in recommendation and (2) the high inference latency caused by generating CoT reasoning. In this work, we explore an alternative approach that shifts from explicit CoT reasoning to compact, information-dense latent reasoning. This approach eliminates the need for explicit CoT generation and improves inference efficiency, as a small set of latent tokens can effectively capture the entire reasoning process. Building on this idea, we propose Reinforced Latent Reasoning for Recommendation (LatentR3), a novel end-to-end training framework that leverages reinforcement learning (RL) to optimize latent reasoning without relying on any CoT data. LatentR3 adopts a two-stage training strategy: first, supervised fine-tuning to initialize the latent reasoning module, followed by pure RL training to encourage exploration through a rule-based reward design. Our RL implementation is based on a modified GRPO algorithm, which reduces computational overhead during training and introduces continuous reward signals for more efficient learning. Extensive experiments demonstrate that LatentR3 enables effective latent reasoning without any direct supervision of the reasoning process, significantly improving performance when integrated with different LLM-based recommendation methods. Our codes are available at https://anonymous.4open.science/r/R3-A278/.
@@ 1 Introduction
Enhancing the reasoning capabilities of Large Language Models (LLMs) has been a central research objective since the LLMs’ emergence, with recent advances—such as DeepSeek-R1 (Guo et al., 2025) and OpenAI-o1 (OpenAI, 2024)—fueling a surge of interest in this direction. By being trained or architected to reason more deliberately through techniques like chain-of-thought (CoT) (Wei et al., 2022) reasoning, LLMs have demonstrated remarkable progress in tackling complex real-world problems (Xu et al., 2025; Shao et al., 2024), rivaling or even surpassing human PhD-level performance in certain cases (OpenAI, 2024). These developments have generated growing enthusiasm for applying LLM reasoning to downstream tasks, including one key AI application for information access—recommender systems (Tsai et al., 2024; Bismay et al., 2024). For recommendation, the core lies in reasoning user preferences from historical behaviors (Tsai et al., 2024), which only implicitly indicate preference. This is well aligned with the strengths of LLM reasoning, offering the potential to unlock new paradigms for personalized recommendation.
@@
Existing approaches to LLM reasoning for recommendation typically rely on explicit textual chain-of-thought (CoT) data to fine-tune models and enhance their reasoning capabilities (Bismay et al., 2024; Tsai et al., 2024; Fang et al., 2025), mirroring trends in general LLM reasoning tasks. However, applying this paradigm to recommendation introduces fundamental challenges at both the tuning and inference stages. First, collecting high-quality CoT data for tuning is inherently difficult: (1) user feedback in recommendation is usually limited to final outcomes, with no access to the underlying reasoning process; and (2) the subjective, personalized nature of preferences (Tsai et al., 2024) makes it both costly and unreliable to manually annotate or synthesize CoT data aligned with individual users. Second, at inference time, these methods require generating explicit CoT reasoning, which is often verbose and incurs substantial latency—an important concern in real-world deployment. These limitations restrict the practical applicability of current methods. To address this, we pose a central question: Can we eliminate the need for explicit CoT reasoning at both the tuning and inference, while still unleashing the reasoning potential of LLM for recommendation?
@@
A promising approach to achieve this is to shift from natural language reasoning to latent reasoning (Hao et al., 2024), where LLMs perform reasoning directly within their hidden representation space, thereby removing the need for explicit textual CoT reasoning. Moreover, hidden states offer much higher information density than textual tokens, allowing compact latent representations to encode complex reasoning processes, thereby reducing the need for lengthy reasoning chains (Su et al., 2025). This would enable more efficient inference. Despite their great potential, existing latent reasoning methods in general domains are not directly applicable to our goal, as they still rely on explicit CoT supervision to learn latent reasoning (Hao et al., 2024), such as by distilling CoT reasoning into latent reasoning.
@@
  • We conduct extensive experiments on real-world datasets, demonstrating that without CoT data, our latent reasoning approach can yield significant performance improvements.
@@ 2 Problem Definition
Let $\mathcal{D}$ denote the collected recommendation data, and let $(u,h,y)\in\mathcal{D}$ denote an instance, where $u$ is a user, $h$ is the user’s historical interactions, and $y$ is the next item the user interacts with. Both $h$ and $y$ are described using textual information (e.g., item titles). To leverage the capabilities of LLMs, we reformulate the recommendation task as a natural language problem. Specifically, for each data point $(u,h,y)$, we convert the historical interactions into a textual prompt $x$, which is then input to the LLM to generate a next-item recommendation. We can also represent a data point by $(x,y)$. Given that user preferences are implicitly and intricately embedded in the historical data, we avoid asking the LLM to produce the recommendation directly. Instead, we introduce a slow thinking process that encourages the model to first perform intermediate reasoning before generating the final output. This can be formalized as:
@@ 3 Methodology
Figure 1: An illustration of the model architecture and training strategy of the proposed LatentR3. On the architectural side, a LatentRATT layer is introduced to generate latent reasoning continuous tokens. The training follows a two-stage framework: the first stage performs warm-up tuning via supervised fine-tuning (SFT), while the second stage applies RL based on a modified GRPO to further optimize the reasoning.
We propose achieving inference-efficient, explicit CoT-free reasoning through reasoning in the latent space and leveraging reinforcement learning to enable reasoning learning. This forms our method, Reinforced Latent Reasoning for Recommendation (LatentR3). As shown in Figure 1, LatentR3 has key designs in both the reasoning architecture and the learning strategy side:
  • Latent Reasoning Architecture. To enable latent reasoning, LatentR3 adds an attention layer on top of the LLM’s final decoding layer, denoted by “LatentRATT". This layer extracts information from the final hidden states of the LLM to generate a latent reasoning token aligned with the LLM’s input embedding space, in an autoregressive manner. Together, the LLM and this generation layer constitute the full reasoning model.
@@ 3.1 Latent Reasoning Architecture
Before generating the final answer, we guide the LLM to first produce reasoning tokens in a continuous latent space to simulate slow thinking. Unlike prior work that directly uses hidden states from the LLM’s decoding layers as latent thoughts within the input embedding space, we introduce an additional attention layer on top of the final decoding layer to explicitly generate latent thought tokens. This layer serves two key purposes: (1) it functions as a dedicated reasoning token generator that aggregates contextual information to produce coherent latent thoughts, and (2) it ensures that the generated tokens are aligned with the LLM’s input embedding space.
@@
$H_{i-1}=LLM^{-1}(x,r_{1},...,r_{i-1}),$ (2)
$r_{i}=\text{LatentRATT}(H_{i-1})[-1],$ (3)
where $LLM^{-1}(\cdot)$ denotes the output of the final decoding layer of the LLM, $r_{1},\dots,r_{i-1}$ are the previously generated latent reasoning tokens, and $H_{i-1}$ is the corresponding sequence of hidden states produced by $LLM^{-1}(\cdot)$. “LatentRATT” refers to our additional attention layer for generating latent reasoning tokens. Finally, $r_{i}\in\mathbb{R}^{d}$ denotes the $i$-th generated latent reasoning token, where $d$ is the dimension size of the LLM’s text token embedding.
Final Generation. After generating $N$ latent reasoning tokens to form the final reasoning thought $r=[r_{1},\dots,r_{N}]$, we concatenate $x$ and $r$ as the input to the LLM for next-item prediction.
@@ 3.2 Reinforced Learning
Inspired by DeepSeek-R1, instead of performing RL from scratch, we first use supervised fine-tuning to warm up the model, providing a strong initial model for RL tuning. We then apply the pure RL to explore better reasoning, forming a two-stage tuning strategy. Next, we introduce each of these stages in detail.
@@ 3.2.1 Warm-up Fine-Tuning
$\text{min}_{\Omega}\quad L_{warm}=-\sum_{(x,y)\in\mathcal{D}}\sum_{i=1}^{|y|}logP_{\theta}(y_{i}|x,r,y_{<i}),$ (4)
where $y_{i}$ denotes the $i$-th token of $y$, $|y|$ denotes the total number of tokens in $y$, and $P_{\theta}(y_{i}|x,r,y_{<i})=\text{LLM}(x,r,y_{<i})$ denotes the predicted probability of $y_{i}$ given the previous generations $y_{<i}$, latent reasoning $r$, and $x$; $\theta$ represents the total model parameters including those of the original LLM and our “LatentRATT” layer, and $L_{warm}$ denotes the total loss.
@@ 3.2.2 RL-based Tuning
Although the model may have gained some preliminary latent reasoning capabilities through SFT, its performance remains insufficient, as it might focus solely on data fitting rather than exploring optimal reasoning paths, leading to suboptimal outcomes. We consider leveraging RL training to facilitate exploration of more diverse reasoning paths, using the SFT results from the first stage as the initialization for further learning.
Given the demonstrated success of GRPO in DeepSeek-R1, we adopt it as the foundation for implementing our RL training. Generally, GRPO operates in four steps: (1) sampling a set of outputs from the old policy model; (2) computing a reward for each output; (3) calculating the advantage by comparing the group average reward; and (4) performing policy updates based on the advantage. To better align with our latent reasoning approach and enhance training efficiency, we introduce specific modifications to the first three steps. We now elaborate on these modifications:
$\bullet$ Sampling. Since our latent reasoning operates in a continuous space, it cannot be directly sampled like discrete textual tokens in standard GRPO. To address this, we leverage the Reparameterization Trick (Kingma et al., 2013, 2015), treating the generated latent reasoning vector as the mean of a Gaussian distribution. For each training sample $(x,y)$ with its corresponding latent reasoning vector $r$, we draw $K$ samples of reasoning. Formally, the $k$-th sampled reasoning ($k\neq 0$) is sampled as:
@@
where $\mathcal{N}(0,\sigma^{2})$ is a Gaussian distribution with zero mean and variance $\sigma^{2}$, $\epsilon$ denotes the sampled noise, and $\sigma$ can be thought of as a hyperparameter controlling the noise strength. Additionally, we include the original $r$ as the 00-th sample, i.e., $r^{0}=r$.
@@
Specifically, we construct the LLM input by combining the original prompt, each sampled latent reasoning vector, and the ground-truth answer. We then compute the model’s perplexity for predicting the ground-truth tokens and use its negative value as the reward to evaluate each latent reasoning direction. Formally, the reward for the $k$-th sampled latent reasoning $r^{k}$ is defined as:
$s^{k}=-exp(-\frac{1}{|y|}\sum_{i=1}^{T}log\pi_{\theta}(y_{i}|x,r^{k},y_{<i})).$ (6)
Here, $\pi_{\theta}$ represents the current policy model (i.e., the current reasoning model), and $\pi_{\theta}(y_{i}|x,r^{k},y_{<i})$ denotes the predicted probability for the $i$-th token in $y$. The rewards corresponding to the group of sampled reasoning instances are denoted as $S=[s^{0},\dots,s^{K}]$. Notably, the rewards are continuous rather than discrete.
@@
[…]
$\bullet$ Policy Update. After computing the advantage, we apply the GRPO loss to update our model. The optimization objective is formulated as follows:
$L_{\text{GRPO}}=\sum_{(x,y)\in\mathcal{D}}-\frac{1}{K}\sum_{k=1}^{K}\frac{1}{|y|}\sum_{i=1}^{|y|}\frac{\pi_{\theta}(y_{i}\mid x,r^{k},y_{<i})}{\pi_{\text{old}}(y_{i}\mid x,r^{k},y_{<i})}\cdot A^{k}-\beta\mathbb{D}_{KL}(\pi_{\theta}||\pi_{ref}).$ (8)
Notably, to reduce computational cost, we update only the “LatentATT” layer while keeping the original LLM layers frozen. As a result, the second KL-divergence term, $\mathbb{D}_{KL}(\pi_{\theta}|\pi_{ref})$, becomes zero in the GRPO implementation.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive reasoning capabilities in complex problem-solving tasks, sparking growing interest in their application to preference reasoning in recommendation systems. Existing methods typically rely on fine-tuning with explicit chain-of-thought (CoT) data. However, these methods face significant practical limitations due to (1) the difficulty of obtaining high-quality CoT data in recommendation and (2) the high inference latency caused by generating CoT reasoning. In this work, we explore an alternative approach that shifts from explicit CoT reasoning to compact, information-dense latent reasoning. This approach eliminates the need for explicit CoT generation and improves inference efficiency, as few latent tokens can effectively capture the entire reasoning process. Building on this idea, we propose Reinforced Latent Reasoning for Recommendation (LatentR3), a novel end-to-end training framework that leverages reinforcement learning (RL) to optimize latent reasoning without relying on any CoT data. LatentR3 adopts a two-stage training strategy: first, supervised fine-tuning to initialize the latent reasoning module, followed by pure RL training to encourage exploration through a rule-based reward design. Our RL implementation is based on a modified GRPO algorithm, which reduces computational overhead during training and introduces continuous reward signals for more efficient learning. Extensive experiments demonstrate that LatentR3 enables effective latent reasoning without any direct supervision of the reasoning process, significantly improving performance when integrated with different LLM-based recommendation methods. Our codes are available at https://github.com/xuwenxinedu/R3 .
@@ 1 Introduction
Enhancing the reasoning capabilities of Large Language Models (LLMs) has been a central research objective since the LLMs’ emergence, with recent advances—such as DeepSeek-R1 (Guo et al., 2025) and OpenAI-o1 (OpenAI, 2024)—fueling a surge of interest in this direction. By being trained or architected to reason more deliberately through techniques like chain-of-thought (CoT) (Wei et al., 2022) reasoning, LLMs have demonstrated remarkable progress in tackling complex real-world problems (Xu et al., 2025; Shao et al., 2024), rivaling or even surpassing human PhD-level performance in certain cases (OpenAI, 2024). These developments have generated growing enthusiasm for applying LLM reasoning to downstream tasks, including one key AI application for information access—recommender systems (Tsai et al., 2024; Bismay et al., 2025). For recommendation, the core lies in reasoning user preferences from historical behaviors (Tsai et al., 2024), which only implicitly indicate preference. This is well aligned with the strengths of LLM reasoning, offering the potential to unlock new paradigms for personalized recommendation.
@@
Existing approaches to LLM reasoning for recommendation typically rely on explicit textual reasoning (i.e., chain-of-thought, CoT) data to fine-tune models and enhance their reasoning capabilities (Bismay et al., 2025; Tsai et al., 2024; Fang et al., 2025), following trends in general LLM reasoning tasks. However, applying this paradigm in recommendation presents fundamental challenges. First, these methods necessitate generating explicit CoT reasoning during inference, which would incur prohibitive computational costs and latency—critical concerns for real-world deployment. Second, collecting high-quality supervision CoT data for effective tuning is difficult: 1) user feedback in recommendation is usually limited to final outcomes, with no access to underlying reasoning; and 2) the subjective, personalized nature of preferences (Tsai et al., 2024) makes manual annotation or synthesis to obtain CoT data both costly and unreliable. All these limitations restrict the practical applicability of current explicit COT methods. To address this, we pose a central question: Can we eliminate the need for explicit CoT reasoning at both the tuning and inference, while still unleashing the reasoning potential of LLM for recommendation?
@@
A promising direction is to move from natural language reasoning to latent reasoning (Hao et al., 2024), where LLMs reason directly in their hidden representation space, eliminating the need for explicit textual CoT reasoning. Moreover, because hidden states have much higher information density than textual tokens, compact latent representations can encode complex reasoning processes. This alleviates the need for lengthy reasoning chains (Su et al., 2025) and thus enables more efficient inference. Yet, existing latent reasoning methods in general domains are not directly applicable to our goal, as they still rely on explicit CoT supervision to learn latent reasoning (Hao et al., 2024), such as by distilling CoT reasoning into latent reasoning.
@@
  • We conduct extensive experiments on real-world datasets, demonstrating that without explicit CoT data, our latent reasoning approach can yield significant performance improvements.
@@ 2 Problem Definition
Let $\mathbb{D}$} denote the collected recommendation data, and let $(u,h,y)\in{\mathbb{D}}$ denote an instance, where $u$ is a user, $h$ is the user’s historical interactions, and $y$ is the next item the user interacts with. Both $h$ and $y$ are described using textual information (e.g., item titles). To leverage the capabilities of LLMs, we reformulate the recommendation task as a natural language problem. Specifically, for each data point $(u,h,y)$, we convert the historical interactions into a textual prompt $x$, which is then input to the LLM to generate a next-item recommendation, denoting the token length as $|x|$. Thereafter, we can also represent a data point by $(x,y)$. Given that user preferences are implicitly and intricately embedded in the historical data, we avoid asking the LLM to produce the recommendation directly. Instead, we introduce a thinking process that encourages the model to first perform intermediate reasoning before generating the final output, which can be formulated as:
@@ 3 Methodology
Figure 1: An illustration of the model architecture and training strategy of the proposed LatentR3. On the architectural side, a LatentRATT layer is introduced to generate latent reasoning continuous tokens. The training follows a two-stage framework: the first stage performs warm-up tuning via supervised fine-tuning (SFT), while the second stage applies RL based on a modified GRPO, termed LR-GRPO, to further optimize the reasoning.
We propose achieving inference-efficient, explicit CoT-free reasoning through reasoning in the latent space and leveraging RL to enable reasoning learning. This forms our method, Reinforced Latent Reasoning for Recommendation (LatentR3). As shown in Figure 1, LatentR3 has key designs in both the reasoning architecture and the learning strategy side:
  • Latent Reasoning Architecture. To enable latent reasoning, LatentR3 adds an attention layer on top of the LLM’s final decoding layer, denoted by "LatentRATT". This layer extracts information from the final hidden states of the LLM to generate latent reasoning tokens (hidden states) aligned with the LLM’s input embedding space, in an autoregressive manner. Together, the LLM and this generation layer constitute the full reasoning model.
@@ 3.1 Latent Reasoning Architecture
Before generating the final answer, we guide the LLM to first produce reasoning tokens in a continuous latent space to simulate slow thinking. Unlike prior work that directly uses hidden states from the LLM’s decoding layers as latent thoughts within the input embedding space, we introduce an additional attention layer on top of the final decoding layer to explicitly generate latent thought tokens. This layer serves two key purposes: (1) functioning as a specialized reasoning token generator that aggregates contextual information to produce coherent latent thoughts, and (2) aligning the latent reasoning tokens more seamlessly with the LLM’s input embedding space.
@@
${\bm{H}}_{i-1}=LLM^{-1}(x,r_{1},...,r_{i-1}),$ (2)
$r_{i}=\text{LatentRATT}({\bm{H}}_{i-1})[-1],$ (3)
where $LLM^{-1}(\cdot)$ denotes the output of the LLM’s final decoding layer; $r_{1},\dots,r_{i-1}$ are the previously generated latent reasoning tokens; and ${\bm{H}}_{i-1}\in\mathbb{R}^{(|x|+i-1)\times d}$ represents the corresponding sequence of hidden states produced by $LLM^{-1}(\cdot)$, with dimensionality $d$. "LatentRATT" refers to our additional attention layer for generating latent reasoning tokens. Finally, $r_{i}\in\mathbb{R}^{d}$ denotes the $i$-th generated latent reasoning token.
Final Generation. After generating $N$ latent reasoning tokens to form the final reasoning thought $r=[r_{1},\dots,r_{N}]$, we concatenate $x$ and $r$ as the input to the LLM for next-item prediction. Here, $N$ is a hyperparameter controlling the length of the latent reasoning sequence, and under our framework, a minimal $N$ (e.g., $N=1$) can achieve strong performance.
@@ 3.2 Reinforced Learning
We leverage RL to facilitate the effective learning of latent reasoning. Inspired by DeepSeek-R1, instead of performing RL from scratch, we first use supervised fine-tuning to warm up the model, providing a strong initial model for RL tuning. We then apply the pure RL to explore better reasoning, forming a two-stage tuning strategy. Next, we introduce each of these stages in detail.
@@ 3.2.1 Warm-up Fine-Tuning
$\text{min}_{\Omega}\quad L_{warm}=-\sum_{(x,y)\in{\mathbb{D}}}\sum_{i=1}^{|y|}logP_{\theta}(y_{i}|x,r,y_{<i}),$ (4)
where $y_{i}$ denotes the $i$-th token of $y$, $|y|$ denotes the total number of tokens in $y$, and $P_{\theta}(y_{i}|x,r,y_{<i})=\text{LLM}(x,r,y_{<i})$ denotes the predicted probability of $y_{i}$ given the previous generations $y_{<i}$, latent reasoning $r$, and $x$; $\theta$ represents the total model parameters including those of the original LLM and our "LatentRATT" layer, and $L_{warm}$ denotes the final total loss.
@@ 3.2.2 RL-based Tuning
Although the model may have gained some preliminary latent reasoning capabilities through SFT, its performance remains insufficient, as it might focus solely on data fitting rather than exploring optimal reasoning paths, leading to suboptimal outcomes. To address this, we consider leveraging RL training to facilitate exploration of more diverse reasoning paths, using the SFT results from the first stage as the initialization for further learning.
Given the demonstrated success of GRPO in DeepSeek-R1 (Guo et al., 2025), we adopt it as the foundation for implementing our RL training. Generally, GRPO operates in four steps: (1) sampling a set of textual outputs from the old policy model; (2) computing a binary reward for each output; (3) calculating the advantage by comparing the group average reward; and (4) performing policy updates based on the advantage. To better align with our latent reasoning approach and enhance training efficiency, we introduce specific modifications to the first three steps, resulting in our new GRPO variant, termed LR-GRPO. In short, (1) we sample latent reasoning tokens in continuous space via the reparameterization trick rather than in textual space, (2) employ perplexity-based rewards, and (3) compute advantages relative to the batch-level average reward instead of the group-level to enhance learning under the new reward design. We now elaborate on these modifications:
$\bullet$ Sampling. Since our latent reasoning operates in a continuous space, it cannot be directly sampled like discrete textual tokens in standard GRPO. To address this, we leverage the Reparameterization Trick (Kingma et al., 2013; 2015), treating the generated latent reasoning vector as the mean of a Gaussian distribution. For each training sample $(x,y)$ with its corresponding latent reasoning vector $r$, we draw $K$ samples of reasoning. Formally, the $k$-th sampled reasoning ($k\neq 1$) is sampled as:
@@
where $\mathcal{N}(0,\sigma^{2})$ is a Gaussian distribution with zero mean and variance $\sigma^{2}$, $\epsilon$ denotes the sampled noise, and $\sigma$ can be thought of as a hyperparameter controlling the noise strength. Additionally, we include the original $r$ as the $1$-st sample, i.e., $r^{1}=r$.
@@
Specifically, we construct the LLM input by combining the original prompt, each sampled latent reasoning vector, and the ground-truth answer. We then compute the model’s perplexity for predicting the ground-truth tokens and use its negative value as the reward to evaluate the quality of each latent reasoning. Formally, the reward for the $k$-th sampled latent reasoning $r^{k}$ is defined as:
$s^{k}=-exp(-\frac{1}{|y|}\sum_{i=1}^{|y|}log\pi_{\theta}(y_{i}|x,r^{k},y_{<i})).$ (6)
Here, $\pi_{\theta}$ represents the current policy model (i.e., the current reasoning model), and $\pi_{\theta}(y_{i}|x,r^{k},y_{<i})$ denotes the predicted probability for the $i$-th token in $y$. The rewards corresponding to the group of sampled reasoning instances are denoted as $S=[s^{1},\dots,s^{K}]$. Notably, the rewards are continuous rather than discrete.
@@
where $\mathbf{S}_{batch}$ denotes the batch of sampling groups, and $\bar{s}_{batch}$ represents the batch-level average reward, computed over the $1$-st sample (i.e., the original reasoning before adding noise) from each group. Formally, it is given by $\bar{s}_{batch}=\frac{1}{N_{group}}\sum_{S=[s^{1},\dots,s^{K}]\in\mathbf{S}_{batch}}s^{1}$, where $N_{group}$ denotes the number of groups in this batch. Lastly, $\|\cdot\|$ denotes the L2 norm.
$\bullet$ Policy Update. After obtaining the advantage, we update our model in a way similar to GRPO. The optimization objective is formulated as follows:
$L_{\text{GRPO}}=\sum_{(x,y)\in\mathcal{D}}-\frac{1}{K}\sum_{k=1}^{K}\frac{1}{|y|}\sum_{i=1}^{|y|}\frac{\pi_{\theta}(y_{i}\mid x,r^{k},y_{<i})}{\pi_{\text{old}}(y_{i}\mid x,r^{k},y_{<i})}\cdot A^{k}-\beta\mathbb{D}_{KL}(\pi_{\theta}||\pi_{ref}).\$(8)
Notably, to reduce computational cost, we update only the "LatentRATT" layer while keeping the original LLM layers frozen. As a result, the second KL-divergence term, $\mathbb{D}_{KL}(\pi_{\theta}|\pi_{ref})$, becomes zero in the GRPO implementation.